OpenStudy (anonymous):

A right circular cylinder with the height of T. The radius of the base is the square root of Q?

OpenStudy (anonymous):

Sounds good. I'll take it!

OpenStudy (anonymous):

can u show me the equation

OpenStudy (anonymous):

The equation for what?

OpenStudy (anonymous):

of the question(up there)

myininaya (myininaya):

do you want volume of the cylinder?

OpenStudy (anonymous):

You've just described a cylinder. What do you want to know about it?

myininaya (myininaya):

surface area?

OpenStudy (anonymous):

umm either one of them

myininaya (myininaya):

Volume=pi*r^2*h=pi*(sqrt(x))^2(T)=pi*x*T

OpenStudy (anonymous):

umm i am a little confuse about this

myininaya (myininaya):

x=Q sorry

OpenStudy (anonymous):

Volume of a cylinder \( = Area_{base} * Height\) Area of a circle \(=\pi r^2\)

OpenStudy (anonymous):

So if Height = T and r = Q. What is the Volume of your cylinder?

OpenStudy (anonymous):

I don't how to do this can u show the step so I can know what the volume of this

OpenStudy (anonymous):

The volume of a cylinder is the (Area of the base) times (the height). The base is a circle. The area of a circle = \(\pi * (radius^2)\). \(\pi \approx 3.14159\) The radius of the base of your cylinder = \(\sqrt{Q}\). The height of your cylinder = T. Using the statements above you can easily compute the volume of your cylinder. Please study what I've said carefully and see if you can figure out how to do this.

OpenStudy (anonymous):

so i use \[\pi \times \sqrt{Q}\]^2 then what next

OpenStudy (anonymous):

Simplify that.

OpenStudy (anonymous):

\(\pi * \sqrt{Q}^2 = \pi * Q\)

OpenStudy (anonymous):

What is that value in relation to your cylinder?

OpenStudy (anonymous):

what that mean

OpenStudy (anonymous):

You use \(\pi * Q\) for what?

OpenStudy (anonymous):

Qpi